NOMINAL GDP GDP DEFLATOR

YEAR (IN BILLIONS) (BASE YEAR 1992)

1993 $725 101.2

1994 $762 102.4

a. What was the growth rate of nominal income between 1993 and 1994? (Note: The growth rate is the percentage change from one period to the next.)

b. What was the growth rate of the GDP deflator between 1993 and 1994?

c. What was real income in 1993 measured in 1992 prices?

d. What was real income in 1994 measured in 1992 prices?

e. What was the growth rate of real income between 1993 and 1994?

f. Was the growth rate of nominal income higher or lower than the growth rate of real income? Explain.
1
Expert's answer
2020-02-18T09:22:22-0500

a) ((762-725)/725)*100=5.1%

b) ((102.4-101.2)/101.2)*100=1.19%

c) (725/101.2)*100=716.4

d) (762/102.4)*100=744.14

e) (744.14/716.4)716.4)*100=3.87%

f) The growth rate of nominal income higher than the growth rate of real income (5.1>3.87), because GDP deflator is higher in 1994 (the real income decreases in a greater extent)